OpenStudy (anonymous):

Which steps transform the graph of y=x^2 to y=-2(x-2)^2+2? A) translate 2 units to the left, translate down 2 units, stretch by the factor 2 😎 translate 2 units to the right, translate up 2 units, stretch by the factor 2 C) reflect across the x-axis, translate 2 units to the left, translate down 2 units, stretch by the factor 2 D) reflect across the x-axis, translate 2 units to the right, translate up 2 units, stretch by the factor 2

OpenStudy (alekos):

y=(x-2)^2 moves the graph to the right by two units y=x^2+2 moves it up 2 y=-x^2 reflects across x-axis y=2x^2 stretches by a factor of 2